### v1

# Contextualizing Recommendation Explanations with LLMs: A User Study

Abstract
Large language models (LLMs) are increasingly prevalent in recommender systems, where LLMs can be used to generate personalized recommendations. Here, we examine how different LLM-generated explanations for movie recommendations affect users’ perceptions of cognitive, affective, and utilitarian needs and consumption intentions. In a pre-registered, between-subject online experiment (N=759) and follow-up interviews (N=30), we compare (a) LLM-generated generic explanations, and (b) LLM-generated contextualized explanations. Our findings show that contextualized explanations (i.e., explanations that incorporate users’ past behaviors) effectively meet users’ cognitive needs while increasing users’ intentions to watch recommended movies. However, adding explanations offers limited benefits in meeting users’ utilitarian and affective needs, raising concerns about the proper design and implications of LLM-generated explanations. Qualitative insights from interviews reveal that referencing users’ past preferences enhances trust and understanding but can feel excessive if overused. Furthermore, users with more active and positive engagement with the recommender system and movie-watching get substantial gains from contextualized explanations. Overall, our research clarifies how LLM-generated recommendations influence users’ motivations and behaviors, providing valuable insights for the future development of user-centric recommender systems, a key element in social media platforms and online ecosystems.
Codehttps://anonymous.4open.science/r/Contextualizing-Recommendation-Explanations-With-LLMs-55D3
1 Introduction
Recommender systems influence human decision-making in daily life, such as purchasing products on e-commerce platforms and choosing movies on streaming services (Angelov et al. 2021; Goyani and Chaurasiya 2020). However, users often fail to understand why they receive specific recommendations. On streaming platforms like Netflix and Disney+, users are presented with recommended movies without any explanation of why those movies might be of interest. Without such explanations, users may find it difficult to judge whether the recommendation is helpful and decide on their consumption. Explanations help users make choices that reflect their preferences and improve decision quality (Vultureanu-Albisi and Badica 2021). As a result, there is a growing research interest in offering explanations behind recommendations (Ko et al. 2022; Zhang and Chen 2020).
The emergence of large language models (LLMs) presents potential for improving these recommendation explanations (Wu et al. 2024; Hou, Leach, and Huang 2024). Some studies have adopted LLMs to produce more natural and human-like explanations. However, current efforts often rely on generic explanation templates (Li, Zhang, and Chen 2020). These generic explanations overlook users’ contexts or preferences, thereby limiting the benefits of the explanations and underutilizing the LLMs’ reasoning capabilities. Existing research seldom provides a comprehensive understanding of how users perceive the received recommendations. Current work also rarely examines whether positive user perceptions translate into users’ consumption intentions.
Motivated by this gap, our study adopts established explanation measurement standards (Balog and Radlinski 2020) and interprets user responses through the lens of the Uses and Gratifications (U&G) framework (Katz, Blumler, and Gurevitch 1973; Lin et al. 2013). The U&G framework categorizes users’ diverse needs, including the desire for understanding and knowledge acquisition (Cognitive needs), the need for emotional connection and engagement (Affective needs), and the focus on practicality and task-oriented goals (Utilitarian needs). This framework provides a valuable perspective to understand how users actively select, consume, and engage with the received information. Our goal is to examine how different LLM-generated recommendation explanation types resonate with specific user needs and, in turn, influence their consumption intentions. We design the following research question:
How do different LLM-generated explanations respond to users’ diverse motivational needs (cognitive, affective, utilitarian) and affect their consumption intentions?
Figure 1: Overview of the study process.
In this study, we conducted a pre-registered, between-subject online experiment involving $N=759$ subjects from the U.S. in the context of movie recommendations. [^1] The experimental protocol was pre-registered at OSF: https://osf.io/j46vx/?view_only=dc2b1f641b33407ab8168736a783055a By drawing upon the U&G framework, we analyze users’ perceptions and intentions towards three types of recommendations generated by GPT-4 Turbo: (1) recommendations with no explanations (NE), (2) recommendations with LLM-generated generic explanations (GE), and (3) recommendations with LLM-generated contextualized explanations (CE). We also drew upon user profile information and conducted follow-up interviews ($N=30$) to interpret our results more deeply. To ensure the robustness of our study design and the generalizability of the results, we conducted supplementary experiments by using a different LLM (Claude 3.5 Sonnet) with an additional 140 users from the U.S. Figure 1 provides an overview of our study flow.
Contributions: This paper demonstrates how LLM-generated, contextualized explanations can more effectively address users’ diverse motivations and increase their consumption intentions. By integrating the U&G framework, our study shows how different explanations affect users’ cognitive, affective, and utilitarian needs. Moreover, we highlight the amplifying role of user engagement with the recommender system, revealing that users who already find the system beneficial gain the most from explanations. These findings bridge advancements in LLMs with user behavior and provide actionable insights to enhance user-centric recommender systems in social media contexts.
2 Related Work
2.1 Explanations in Recommender Systems
In social media and online communities, recommendations help users to find items aligned with their preferences. Initial research centered on improving accuracy (Ma et al. 2023; Wu et al. 2023), but recent work emphasizes explainability to foster trust, transparency, and user satisfaction (Burkart and Huber 2021). Early approaches integrated explainability into model architectures, such as explainable matrix factorization leveraging user similarities (Abdollahi and Nasraoui 2017), tree-enhanced embeddings (Wang et al. 2018), and attention-based neural models (Seo et al. 2017; Chen et al. 2018). More recent efforts have expanded explanation strategies, ranging from visual (Tsai and Brusilovsky 2019) and comparative methods (Le and Lauw 2021) to sentiment-oriented narratives (Chen, Yan, and Wang 2019), as well as traditional item-, content-, and popularity-based explanations (Sarwar et al. 2001; Pazzani and Billsus 2007; Yang 2016). These approaches highlight the importance of making recommendations more understandable, interactive, and socially aligned to support user trust and transparency across online communities.
Beyond these approaches, LLMs offer new opportunities for generating better explanations within the dynamic, user-driven landscapes (Mohanty 2024). By producing natural language explanations, LLMs can improve user satisfaction and trust, especially in domains where social context and user interaction are central (Chang, Harper, and Terveen 2016). For instance, LLMs have been used to create more detailed item descriptions (Acharya, Singh, and Onoe 2023) or to enable ChatGPT-based recommenders to generate user-friendly explanations (Deldjoo 2024). Recent research suggests that more contextualized, user-centered explanations could enhance trust and engagement (Silva et al. 2024; Lyu et al. 2024), reinforcing the value of LLM-driven strategies that align closely with the evolving social and informational needs of online communities.
2.2 User Perceptions of Recommendation Explanations
Users increasingly rely on recommender systems to navigate overwhelming volumes of information. Research grounded in the U&G framework highlights that individuals engage with media to fulfill specific needs, which can be categorized as affective, cognitive, and utilitarian motivations (Katz, Blumler, and Gurevitch 1973; Korhan and Ersoy 2016). These motivations directly influence consumption intentions, as users are more likely to engage with media or systems that effectively address their various needs. High-quality explanations within recommender systems can significantly influence user engagement, satisfaction, and decision-making, particularly in socially dynamic environments where these systems satisfy informational and social needs (Mandl et al. 2011; Jameson et al. 2015). For instance, explanations can enhance perceived usefulness and trust, aligning with motivation needs such as ease of use and task efficiency (Korhan and Ersoy 2016). Extant research also offers valuable insights into factors shaping user perceptions. For example, evidence shows that users may prefer human-generated recommendations, particularly for hedonic products (Wien and Peluso 2021). Similarly, explanations reflecting user feedback and personalization resonate with individuals who seek tailored, value-reinforcing content (Lu et al. 2023). Differences in user profiles further emphasize the importance of detailed and specific explanations, particularly for users with high cognitive engagement who value transparency and trust (Chatti et al. 2022).
Natural language explanations further enable systems to fulfill diverse needs, such as reducing cognitive effort, fostering trust, and enhancing community integration (Chang, Harper, and Terveen 2016). In parallel, researchers have developed evaluation metrics to address these user-centric demands beyond traditional measures like BLEU and ROUGE (Papineni et al. 2002; Lin 2004). User-centered metrics now consider dimensions such as effectiveness, persuasiveness, satisfaction, and trust, emphasizing user needs in complex, community-driven contexts (Zhou et al. 2021; Ehsan et al. 2019; Hoffman et al. 2023).
2.3 Research Gap
While the importance of recommendation explanations is widely acknowledged, research seldom examines how users interpret and respond to them. Furthermore, extant studies often rely on generic templates and have not fully explored the potential of LLM-generated contextualized explanations to address users’ cognitive, affective, and utilitarian needs (see Table 1).
To address this gap, our study provides the first comprehensive user evaluation of LLM-generated contextualized explanations through the lens of the U&G framework. By examining how these explanations resonate with users’ multifaceted needs and influence their intentions, we aim to offer deeper insights into designing recommendation explanations that are more aligned with user motivations.
Kunkel et al. Silva et al. Lubos et al. Our study Contextualized explanations Generic explanations Cognitive needs Affective needs Utilitarian needs Consumption intentions
Table 1: Comparison of research focus between previous studies and our work.
3 Method
3.1 Overview
We conducted an online between-subject experiment, where subjects ($N=759$) were randomly assigned to one of three experimental conditions: (1) recommendations with no explanations (NE), (2) recommendations with LLM-generated generic explanations (GE), and (3) recommendations with LLM-generated contextualized explanations (CE).
In the experiment, users first selected movies they had previously watched from a given set to establish the viewing history, based on which LLMs generated movie recommendations and explanations depending on the experimental condition. [^2] We used GPT-4 Turbo (version: gpt-4-1106-preview), which has knowledge of world events up to April 2023. For the experiment in the supplement, we used Claude 3.5 Sonnet. After receiving the corresponding explanation, users provided a series of ratings reflecting how well the explanation met underlying motivational needs, as well as their resulting behavioral intentions:
  • Cognitive needs: reflecting users’ desire for understanding and credibility in the recommendation process, assessed through Scrutability, Transparency, and Trust.
  • Affective needs: addressing the emotional and motivational dimensions of the user experience, evaluated through Persuasiveness and Satisfaction.
  • Utilitarian needs: focusing on practical, goal-oriented aspects of the recommendation, captured by Effectiveness, Efficiency, and Helpfulness.
  • Consumption intentions: reflect users’ behavior outcome, captured by UserIntent.
Dimensions composing the needs and intentions are established for assessing explanations in prior user studies (Balog and Radlinski 2020). We organize the dimensions under the U&G framework to depict users’ active pursuit of various needs from explanations. Given the statement for each dimension, users rated their agreement on a 7-point Likert scale from 1 (“Extremely Disagree”) to 7 (“Extremely Agree”). The higher-level intentions and needs are averaged on the specific dimensions (e.g., Cognitive is the average of Scrutability, Transparency, and Trust). A summary of the main variables and survey questions is in Table 2. After the survey, we randomly invited users to conduct follow-up interviews ($N=30$) to see their latent motivations across different explanations.
Variable Questions
Cognitive Needs
Transparency The text helps me to understand what the recommendation is based on.
Trust The text helps me to trust the recommendation.
Scrutability The text would allow me to give feedback on how well my preferences have been understood.
Affective Needs
Persuasiveness The text makes me want to watch this movie.
Satisfaction The text would improve how easy it is to pick a recommendation.
Utilirian Needs
Effectiveness The text helps me to determine how well I will like this movie.
Efficiency The text helps me to decide faster if I will like this movie.
Helpfulness How helpful do you find the recommendation?
Behvaiour Outcome
UserIntent How willing are you to watch this recommended movie?
Table 2: Summary of main variables and survey questions.
3.2 Subjects
Subjects were recruited from the U.S. via Cint. [^3] Cint: https://www.cint.com/ The online platform offers a built-in functionality for stratified sampling of subjects so that the sample is representative of the U.S. population. Subjects are remunerated with a fixed amount of around \$15 per hour.
As preregistered, we conducted a pilot study involving 50 subjects to assess the survey’s functionality and design. Informed by other user studies (Kang 2021; Lakens 2022), we performed a power analysis using G*Power 3.1 software (Faul et al. 2007, 2009). For multiple comparisons, we selected $F$-tests and chose a one-way ANOVA design with fixed effects (Lee and Lee 2018). We set the significance level ($\alpha$) to 0.05 and the power ($1-\beta$) to 0.95, with the adjusted effect size set at 0.1. The power analysis based on the pilot study results suggested that we need at least 750 subjects for the experiment.
Initially, 1401 subjects completed the survey. We discarded responses from subjects failing an initial attention check. Subsequently, following prior research (Pennycook, Cannon, and Rand 2018) and our pre-registration, a subset of subjects were excluded for three reasons: (i) subjects who self-reported to have answered randomly, (ii) failed the second attention check, (iii) responded excessively fast (i.e., the duration was less than around 6 seconds for responding to the user perceptions). These users still received remuneration after the exclusion. Eventually, due to our strict exclusion criteria, 759 of the 1,401 subjects qualified, exceeding our estimated sample size of 750 to ensure the results achieve the desired statistical power.
Among the 759 subjects, 387 are women, 370 are men, and two are non-binary. The sample has a mean age of 52.85 (SD = 16.01) and a range of 70 years (18 to 88). Most subjects have at least a high school diploma or equivalent.
3.3 Materials
We selected GPT-4 Turbo to generate movie recommendations for the main experiment. We conducted an additional experiment with Claude 3.5 Sonnet to validate the robustness of our study design and the generalizability of our findings (see supplementary materials). We precompiled all recommendations and explanations in a corpus to ensure that there was no variability in the outcomes across different LLM calls.
To create the corpus, we first selected 10 movies with diverse topics from the IMDb Top 250 list [^4] https://www.imdb.com/chart/top/ to compile a movie shortlist $\mathcal{M}$. These movies were chosen based on their high ratings or popularity. We ensured that the shortlist had broad coverage of user viewing preferences. Subsequently, we implemented a sequential recommendation strategy following the algorithm in Table 5 in Appendix A. With this strategy, sujects with the same viewing history and assigned conditions received the identical recommendation..
We followed previous practices (Giray 2023) when designing the prompts. Table 3 shows examples of contextualized and generic explanations. We can see a clear difference: while both explanations incorporate various elements (e.g., feature, popularity, similarity) to support the recommendations, the contextualized explanations naturally integrate the recommended movies with the user’s viewing history, explicitly highlighting how the user’s inputs shape the recommendation. Meanwhile, the generic explanations focus exclusively on the recommended movies. All generated movies were cross-checked with the IMDB dataset [^5] https://developer.imdb.com/non-commercial-datasets/ to ensure no hallucinations from the LLMs. Additionally, the corpus was validated by two experts with strong expertise in the field to ensure that all explanations used in the experiment were consistent and of fair quality.
System Prompt You are a movie recommendation assistant. Aim for clear, straightforward language that a broad audience can easily understand. Output each recommendation into three segments: The movie’s details, formatted as “Title - Release Year - Director”; A concise summary within 40 words that outlines the storyline; An explanation within 100 words. The explanation should focus on concrete aspects such as specific character arcs, notable plot developments, or memorable scenes.
User Prompt - CE For {Movie}, provide one recommendation for a user with a viewing history: {User viewing history}. Highlight how it directly appeals to the provided viewing history.
User Prompt - GE For {Movie}, write a recommendation explanation.
Table 3: Examples of contextualized and generic explanations generated by GPT-4 Turbo.
3.4 Procedure
During the experiment, we collected users’ viewing history and assigned them to one of our three conditions. Based on the condition assignment and viewing history, we generated the LLM-generated recommendations for them and collected their evaluations. The entire procedure was conducted online, consisting of the following steps:
Step 1. Users visit a welcome page where they provide informed consent to proceed with the experiment.
Step 2. Users read a short paragraph about the movie recommendation and respond to an attention check. Users who fail to pass are re-directed to the end of the survey. The attention check also serves to assess English proficiency.
Step 3. Users read about the purpose and functionality of a recommender system.
Step 4. Users select movies they have seen from the movie shortlist $\mathcal{M}$.
Step 5. Users provide the frequency of their movie consumption, their usage frequency, and the perceived helpfulness of recommender systems.
Step 6. Users are randomly assigned to one of our three conditions and shown corresponding recommendations.
Step 7. Following each recommendation, users express their perceptions across the dimensions of cognitive, affective, and utilitarian needs on a 7-point Likert scale. Users also state the recommendation’s helpfulness and express their intentions to consume the movie. Steps 6 and 7 are repeated for all movies selected by users from $\mathcal{M}$.
Step 8. Users go through a second set of attention checks.
Step 9. Users provide demographic information, i.e., age, gender, and education.
Step 10. Users participate in an honesty test where they are asked whether they answered all questions honestly, while being guaranteed that compensation would not be contingent on their answers.
Step 11. Users receive a debrief of the experiment and are informed that LLMs power all movie recommendations.
Afterward, we randomly reached out to $N=30$ users for the follow-up interviews. [^6] Detailed interview questions in Appendix B. We integrated their responses to help interpret the quantitative analysis results and investigate how explanations affect user motivational needs.
3.5 Analyses
We conducted analyses based on user perceptions in the survey. Each participant could provide up to 10 ratings, resulting in a total of 2,732 ratings from 759 participants in the main experiment (916 in NE, 898 in GE, and 918 in CE). We used hypothesis testing to identify differences in perceptions and consumption intentions across recommendation types, followed by mixed-linear regression to examine how user characteristics explain variation in both variables. Additionally, we conducted follow-up interviews to further interpret the quantitative results.
Hypothesis testing: We hypothesized that there are significant differences in the perceptions and consumption intentions between users receiving different recommendations. We conducted hypothesis testing on the four dependent variables of interest: Cognitive, Affective, Utilitarian, and UserIntent across the different conditions. Since the data did not meet the normality assumption (Lilliefors 1967), we used a Mann–Whitney test with Holm correction to identify pairwise differences between conditions (McHugh 2011). Later, we use $\mu$ to report the mean.
Mixed-effects linear regression. We used linear mixed-effects regression models to check further if any factors explained the variations in the three variables of interest. [^7] We also calculated the variance inflation factor (VIF) for all independent variables. All VIF values were below 5, indicating that multicollinearity should not bias the estimates (Akinwande, Dikko, and Samson 2015). We set NE as the reference condition so that our Treatment variables represent CE or GE. Given that each subject provides multiple ratings, we further included subject-specific random effects to control for between-subject heterogeneity (Linden and Hönekopp 2021). Additionally, to control for the habituation effect caused by repetitive measurements (Gaito 1961; Jankowski 2021), we included a control variable for the recommendation display order (i.e., DisplayOrder). Our final regression model is:
$\emph{Y}_{ij}$ $=\beta_{0}+\beta_{1}\emph{Treatment}_{j}+\beta_{2}\emph{Age}_{j}+\beta_{3}\emph{Gender}_{j}+\beta_{4}\emph{Education}_{j}$
$\quad+\beta_{5}\emph{MovieConsumption}_{j}+\beta_{6}\emph{RecSysFrequency}_{j}$ (1)
$\quad+\beta_{7}\emph{RecSysUtility}_{j}+\beta_{8}\emph{DisplayOrder}_{j}+u_{0j}+\epsilon_{ij}$
where ${Y}_{ij}$ denotes the dependent variables (i.e., Cognitive, Affective, Utilitarian, and UserIntent) for the $i$-th observation of the $j$-th subject, with intercepts $\beta_{0}$, subject-level random effects $u_{0j}$, coefficients $\beta_{1}$ through $\beta_{8}$, and an error term $\epsilon_{ij}$.
3.6 Ethical Considerations
We respect the privacy and agency of all subjects potentially impacted by this work and take steps to protect their privacy. The experiment design has obtained approval from the Ethics Commission at the University of Lausanne. All data were collected anonymously, and any personally identifiable features were removed. All steps during data collection and analyses followed standards for ethical research (Rivers and Lewis 2014). We also informed users that movie recommendations and explanations were generated by an LLM.
4 Results
Below, we state the results of the hypothesis tests. The regression results are also shown in Table 4.
4.1 Cognitive Needs
We observed a significant difference (Holm-adjusted $p<0.05$) in the distribution of the aggregated Cognitive measure (i.e., Scrutability, Transparency, Trust) between NE ($\mu=4.87$, $95\%$ CI = [4.78, 4.95]) and CE ($\mu=5.22$, $95\%$ CI = [5.14, 5.30]). We also found a significant difference between GE and CE (Holm-adjusted $p<0.05$), with CE surpassing GE ($\mu=4.90$, $95\%$ CI = [4.81, 4.99]). However, there was no statistical difference between NE and GE (Holm-adjusted $p>0.05$), suggesting that generic explanations do not differ markedly from no explanations in meeting users’ cognitive needs.
In the mixed-effects regression, both RecSysUtility ($\beta=0.319$, $p<0.05$) and MovieConsumption ($\beta=0.149$, $p<0.05$) demonstrated positive, significant effects on Cognitive, indicating that participants who perceive the recommender system as more useful and who watch movies more frequently tend to rate higher in Cognitive.
In the follow-up interviews, several participants expressed increased trust and clarity when receiving the contextualized explanations. One participant noted: “When it told me the rationale for suggesting the movie, I could decide whether or not to watch it.” Another participant stated: “Including the movie I had watched and comparing the similarities made me trust the recommendation more.” Interestingly, some users also noted that references to movies they disliked helped them filter out unwanted content, improving their sense of control and overall transparency.
These quantitative and qualitative findings suggest that contextualized explanations, by explicitly linking a user’s past and present preferences, can bolster trust and understanding in a way that generic or absent explanations do not. The more personal and concrete presentation of information serves users’ cognitive needs by explaining why some recommendations might be a good match, and others might not align with their tastes, ultimately leading to a deeper sense of scrutiny, transparency, and trust.
Figure 2: Cognitive ratings by explanation type (shown variable: Cognitive). Significance levels between conditions are indicated: ${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$.
4.2 Affective Needs
We found no overall difference in the aggregated Affective measure (i.e., Persuasiveness, Satisfaction) among NE ($\mu=4.97$, $95\%$ CI = [4.88, 5.05]), GE ($\mu=4.92$, $95\%$ CI = [4.83, 5.01]), and CE ($\mu=5.02$, $95\%$ CI = [4.93, 5.11]). However, examining each subdimension showed that Persuasiveness differed significantly between GE and CE (Holm-adjusted $p<0.05$), indicating that contextualized explanations can make recommendations more compelling than generic explanations, even if this advantage does not fully carry over to overall perceived affective needs.
The mixed-effects regression highlights two user-engagement variables, RecSysUtility ($\beta=0.311$, $p<0.05$) and MovieConsumption ($\beta=0.167$, $p<0.05$), as positive, significant predictors of affective responses. This suggests that users who view the recommender system as helpful and frequently watch movies are more likely to feel emotionally drawn in, regardless of the explanation style.
Our interviews offer deeper insights into how these affective needs take shape. Several participants emphasized that a good explanation heightened their eagerness to watch the movie: “They made me want to watch them more than I would have before with their explanations.” Others found “overly pushy” or “over the top” descriptions off-putting, highlighting the delicate balance between building excitement and overselling the film. Some appreciated specific references to actors or plot details, while others disliked spoiler-like content, favoring a brief, suggestive teaser. These varying reactions prove that users’ emotional engagement hinges on contextual cues and personal taste, mood, and openness to promotional language.
These findings imply that satisfying affective needs, such as feeling entertained, inspired, or emotionally “in tune”, goes beyond simply adding contextual detail. Users’ enjoyment levels are tightly bound to their existing trust in the system, general interest in watching movies, and thresholds for promotional or spoiler content. Consequently, while contextualized explanations enhance persuasiveness for some users, effectively catering to affective motivations requires a more nuanced approach that recognizes diverse emotional triggers and preferences.
Figure 3: Affective ratings by explanation type (shown variable: Affective).
4.3 Utilitarian Needs
There is no overall difference in the aggregated Utilitarian measure (i.e., Effectiveness, Efficiency, Helpfulness) across NE ($\mu=5.24$, $95\%$ CI = [5.16, 5.32]), GE ($\mu=5.27$, $95\%$ CI = [5.19, 5.36]), and CE ($\mu=5.27$, $95\%$ CI = [5.18, 5.35]). However, a closer examination at Helpfulness revealed that NE ($\mu=5.07$, $95\%$ CI = [4.98, 5.17]) versus CE ($\mu=5.23$, $95\%$ CI = [5.13, 5.33]) was significant (Holm-adjusted $p<0.05$), suggesting that contextualized explanations can offer an advantage over having no explanation in practice by helping users assess whether a movie meets their needs. By contrast, there was no difference between GE and CE in Helpfulness, indicating that once any form of explanation is provided, adding contextual details may not further enhance users’ perceived practical value in selecting movies.
The mixed-effects regression showed RecSysUtility ($\beta=0.292,p<0.05$) and MovieConsumption ($\beta=0.181,p<0.05$) both positively influenced Utilitarian, again implying that users who closely engage with recommender systems and movies, tend to see greater practical value in the recommendations regardless of explanation style. Female participants also reported slightly higher utilitarian scores (versus male, $p<0.05$).
Our interview provides further evidence. Several participants appreciated having clear, concise details about the plot or genre, emphasizing how these insights save time in deciding whether a movie fits their tastes: “The more it detailed the movie, the more it helped me decide.” Some also suggested that referencing both liked and disliked movies aided them in filtering out poor matches and reinforcing good ones: “I was able to assess whether I wanted to watch it, based on what I liked before, and also on what I knew I didn’t enjoy.” Others felt the explanations were too vague or enthusiastic, hindering their ability to gauge true quality.
Although contextualized explanations may slightly outperform no explanations in offering helpful guidance, overall utilitarian gains do not appear large when considering Effectiveness or Efficiency. These findings suggest that users’ utilitarian needs stem less from how the explanation is framed and more from whether the system is already perceived as useful and aligns with existing consumption habits. Nonetheless, highlighting specific points of similarity or dissimilarity can provide an additional benefit in helping users quickly identify which recommendations are worth pursuing and which to skip.
Figure 4: Utilitarian ratings by explanation type (shown variable: Utilitarian).
4.4 Consumption Intentions
Knowing users’ consumption intentions (measured by UserIntent) helps us understand how explanations translate into actual viewing decisions. Pairwise comparisons confirm that CE ($\mu=4.70$, $95\%$ CI = [4.59, 4.81]) significantly differs from GE ($\mu=4.31$, $95\%$ CI = [4.18, 4.43]) (Holm-adjusted $p<0.05$). The contrasts between NE ($\mu=4.53$, $95\%$ CI = [4.42, 4.65]) and CE, as well as NE and GE, are also significant (Holm-adjusted $p<0.05$). This pattern suggests that the existence of explanations and contextualization fosters greater motivation to watch a recommended movie.
Mixed-effects regression indicates that CE increases consumption intentions by 0.240 points on average ($\beta=0.240,p<0.05$) compared to GE. RecSysUtility ($\beta=0.336,p<0.05$) and MovieConsumption ($\beta=0.143,p<0.05$) both increase the likelihood of watching the recommended title. Meanwhile, older participants ($\beta=-0.139,p<0.05$) and female users ($\beta=-0.227,p<0.05$) expressed slightly lower consumption intentions.
The qualitative interview responses underscore how alignment with personal taste can motivate users to watch the recommended movies. Several participants mentioned that recognizable themes, plot twists, or well-known actors greatly influenced their choices. One participant explained: “An interesting twist or variation from an expected plot development intrigues me, as does the presence of an actor I like.” Meanwhile, others highlighted that the mention of disliked or irrelevant elements quickly led them to dismiss a recommendation: “Some movies I simply will not watch … if it’s not my kind of genre, no explanation can convince me.” In some instances, participants weighed social factors such as the opinions of friends and family members, underscoring how shared viewing decisions can override individual preferences if, for example, a partner or group prefers a different type of film.
These results imply that while contextualized explanations provide an edge over generic ones, individual beliefs about system helpfulness and prior consumption habits remain critical in determining whether users actually watch a recommended movie. Explaining “why this movie” helps, but so does the user’s fundamental perception of the recommender’s reliability and personal engagement with the medium. Ultimately, consumption intentions emerge from a synthesis of explanation quality, personal taste, and social or experiential factors, illustrating that users weigh both the persuasive power of contextualization and their unique circumstances.
Figure 5: Consumption intentions by explanation type (shown variable: UserIntent). Significance levels between conditions are indicated: ${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$.
Cognitive Affective Utilitarian UserIntent
Intercept
4.870***
4.906***
5.084***
4.642***
Treatment (GE)
0.030
--0.010
0.082
--0.180
Treatment (CE)
0.366***
0.125
0.118
0.240*
Gender (Female)
0.094
0.134
0.284***
--0.227*
Age
--0.061
--0.076
0.025
--0.139*
Education
0.049
0.047
0.019
0.060
RecSysUtility
0.319***
0.311***
0.292***
0.336***
RecSysFrequency
0.013
--0.004
--0.043
0.106
MovieConsumption
0.149***
0.167***
0.181***
0.143*
DisplaypOrder
--0.018
--0.046
--0.031
--0.022
Significance levels: ${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$
Table 4: Mixed-effect linear regression results. Coefficients are shown with standard errors in parentheses.
5 Supplementary Experiments
To supplement the main experiment, we expanded the experiment to a larger scope by leveraging another LLM (Claude 3.5 Sonnet).
Given that GPT-4 Turbo is one of the most powerful LLMs, we tested whether the superior performance of contextualization could be generalized to other LLMs. We generated another recommendation corpus using Claude 3.5 Sonnet and replicated the experiments on the U.S. samples ($N=140$, 71 are women, 68 are men, and one is non-binary). The results are aligned with our main findings, validating the advantages of LLMs in generating contextualized explanations in terms of meeting users’ motivational needs and increasing their consumption intent. The results and interpretations are provided in Appendix C.
In summary, the supplementary experiment yielded results that supported the main experiment’s findings and conclusions, thereby validating the robustness of our study design and the generalizability of the findings.
6 Discussion
6.1 Implications of Findings
Recommender systems play a central role in shaping user decisions, particularly influential in the context of web and social media. Given the far-reaching implications of these systems, there is growing demand from users and regulators for transparency, emotional connection, or efficiency in decision-making. This highlights the value of high-quality and context-aware recommendation explanations. However, many existing recommender systems do not or fail to provide explanations that address users’ diverse needs (Li, Zhang, and Chen 2020; Lu et al. 2023).
Our study demonstrates that LLM-generated contextualized explanations can overcome these limitations by satisfying users’ cognitive needs, ultimately increasing their consumption intentions. Also, our findings reveal the insufficiency of LLM-generated explanations in meeting users’ utilitarian and affective needs, which raises concerns about the proper design and implications of LLM-generated personalized content.
Benefits of contextualization. Our results show that contextualized explanations significantly satisfy users’ needs and their intention to watch the recommended movies. This is particularly evident in users’ cognitive needs (Scrutability, Transparency, Trust), where contextualization not only promotes a clearer understanding of why specific titles are recommended but also encourages users to scrutinize recommendations more carefully. The increase in trust and perceived transparency shows the importance of tailoring explanations to individual needs. Offering personally relevant and meaningful recommendations can encourage users to act on these suggestions, especially in entertainment and e-commerce settings, where a personalized touch can be key to customer retention.
LLMs in the recommendation. As LLMs evolve, their capacity to generate richer, more personalized explanations will continue to grow (Feuerriegel et al. 2025). This capability is particularly relevant for companies like Amazon and Netflix, where personalization is a competitive differentiator (Burroughs 2019). Our findings highlight how integrating LLMs into recommender systems can satisfy users’ cognitive needs and increase their consumption intention. Beyond driving engagement, context-aware explanations may help companies fulfill emerging regulatory requirements, such as the EU’s AI Act, which calls for transparency and the “right to explanation” in AI-driven decisions (Veale and Borgesius 2021; You and Suh 2024).
System utility and positive feedback loops. Our study suggests that a user’s active and positive engagement with the recommender system amplifies the benefits of contextualization. When participants view the system as helpful and relevant to their interests, context-aware explanations further boost their ratings of recommendation quality, trust, and helpfulness. This synergy can create a positive feedback loop: as users find the system increasingly beneficial, they engage more deeply, and contextualized explanations, in turn, reinforce these favorable perceptions. Moreover, frequent movie watchers also exhibit higher cognitive, affective, and utilitarian evaluations. Increasing user engagement through personalized interfaces, easy-to-use features, or curated content could, therefore, magnify the impact of contextualized explanations (Lacker and Way 2024). Companies seeking to cultivate sustained user relationships can harness this dynamic by combining context-rich recommendation details with user education or incentives that further encourage exploration and use of the system.
Avoiding redundant information. An interesting insight from the experiment is that some participants found overly “pushy” or “over the top” descriptions off-putting. In contrast, others appreciated detailed explanations. Although contextualized explanations can meet various user needs, avoiding irrelevant or excessive information that increases cognitive load is critical. Our findings suggest that, for specific users, no explanation is preferable to offering a generic one. This reflects a balancing act: well-crafted, context-specific explanations can significantly improve user perceptions, but lengthy or poorly targeted explanations can become a distraction.
In industries such as travel and finance, where decisions can be complex, selective use of contextualized explanations could help users make informed choices without overwhelming them. By highlighting only the most pertinent details, companies can streamline the decision-making process, thus promoting higher engagement and satisfaction.
6.2 Limitations
While our study offers valuable insights into user perceptions of LLM-generated explanations for movie recommendations, several limitations should be noted.
First, we primarily recruited English-speaking users. Future research should include more diverse populations to strengthen the generalizability of our findings.
Second, the effectiveness of the explanations may have been influenced by the performance of the underlying LLMs. Although we used state-of-the-art models (GPT-4 Turbo and Claude 3.5 Sonnet), we should notice that future advancements in LLMs could further shape or change user perceptions.
Third, our distinction between contextualized and generic explanations may oversimplify the explanation types. Future studies should explore a broader variety of explanation strategies to better understand their impact in different contexts and on diverse user groups.
Lastly, our experimental design did not fully capture the dynamic real-world user interactions with recommender systems. User perceptions might change after consuming the recommended movies, which is not covered by this study. This highlights the necessity of conducting the user study in real scenarios.
Despite these limitations, our study provides a strong foundation for further research into the role of LLM-generated explanations in recommender systems, emphasizing the importance of contextualization in improving user experience and engagement.
7 Conclusion
Our findings show how and why recommendation explanations meet cognitive, affective, and utilitarian motivations under the U&G perspective. Our study confirms that LLM-generated contextualized explanations effectively address users’ cognitive needs and increase their consumption intentions. Notably, recommender systems should also avoid overwhelming users with irrelevant information that can hinder decision-making or erode trust. Future recommender system designs that integrate contextualized LLM-generated explanations can more effectively serve a wide range of user motivations, ultimately improving user satisfaction and leading to actionable engagement with the recommended content.
Appendix A Recommendation Corpus Generation
Recommendation Corpus Generation
Input: Movie shortlist $\mathcal{M}$
Output: Corpus $\mathcal{C}$ containing recommendations and explanations
1 for each $m_{i}\in\mathcal{M}$ do
   $m_{i}$: Represents a single movie in the shortlist $\mathcal{M}$
2    for $k=1$ to 4 do
   $k$: The recommendation order (1st, 2nd, 3rd, or 4th)
3     if $k=1$ then
4      Generate $r_{i1}$ from $m_{i}$
     $r_{ik}$: The $k$-th recommendation for the $i$-th movie
5     else
6      Generate $r_{ik}$ from $\{m_{i},r_{i1},\dots,r_{i(k-1)}\}$
     Previous recommendations are incorporated to generate $r_{ik}$
7     end if
8     Generate $e_{ikc}$ and $e_{ikg}$ for $r_{ik}$
    $e_{ikc}$: Contextualized Explanation for $r_{ik}$
    $e_{ikg}$: Generic Explanation for $r_{ik}$
9     Save $r_{ik}$, $e_{ikc}$, and $e_{ikg}$ in $\mathcal{C}$
    Corpus $\mathcal{C}$ stores the recommendations and their explanations
10    end for
11 end for
Table 5: Recommendation corpus preparation steps. The corpus $\mathcal{C}$ contains recommendations $r_{ik}$ for each movie $m_{i}\in\mathcal{M}$, with contextualized explanations ($e_{ikc}$) and generic explanations ($e_{ikg}$) for each recommendation.
Appendix B Interview Questions
  • “What made you trust or doubt the recommendation after seeing the explanation?”
  • “How did the explanation affect your overall feeling about the movie? Why did it make you feel that way?”
  • “How did the explanation assist (or fail to assist) you in understanding if the movie fit your interests or needs?”
  • “After reading the explanation, what influenced you most in deciding whether to watch the movie or not?”
  • “How do you think your profile, such as your demographics or viewing habits, influenced how you felt about the explanation you received?”
Appendix C Supplementary Experiment (with Claude 3.5 Sonnet)
We replicated the experiment on a smaller U.S. sample ($N=140$) using Claude 3.5 Sonnet. The results align with our main study, showing that contextualization improves overall explanation quality, particularly in terms of persuasiveness, transparency, and trust. Moreover, contextualized explanations increase users’ consumption intentions. Additionally, the positive effect of RecSysUtility is maintained, even with another underlying LLM instead of GPT-4 Turbo. While some minor discrepancies emerged between this supplementary experiment and the main experiment, the core findings remain consistent. This supports the generalizability of our study’s conclusions. Figure 6, Figure 7, Figure 8, and Figure 9 show the user ratings of motivational needs and consumption intentions in this experiment.
Figure 6: Cognitive ratings by explanation type (shown variable: Cognitive). Significance levels between conditions are indicated: ${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$.
Figure 7: Affective ratings by explanation type (shown variable: Affective).
Figure 8: Utilitarian ratings by explanation type (shown variable: Utilitarian).
Figure 9: Consumption intentions by explanation type (shown variable: UserIntent). Significance levels between conditions are indicated: ${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$.
Cognitive Affective Utilitarian UserIntent
Intercept
4.823***
4.807***
5.103***
4.559***
Treatment (GE)
0.097
0.112
--0.055
0.101
Treatment (CE)
0.509*
0.534*
0.331
0.730*
Gender (Female)
--0.126
--0.228
0.064
--0.499
Age
0.269*
0.234*
0.274*
0.145
Education
--0.006
--0.040
--0.052
0.053
RecSysUtility
0.217*
0.155
0.268
0.418*
RecSysFrequency
0.156
0.146
0.076
0.030
MovieConsumption
0.044
--0.018
--0.045
--0.148
Significance levels: ${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$
Table 6: Mixed-effects linear regression results for the four dependent variables: Cognitive, Affective, Utilitarian, and UserIntent. Coefficients are shown with standard errors in parentheses.

### v2

@@
 Abstract
-Large language models (LLMs) are increasingly prevalent in recommender systems, where LLMs can be used to generate personalized recommendations.
-Here, we examine how different LLM-generated explanations for movie recommendations affect users’ perceptions of cognitive, affective, and utilitarian needs and consumption intentions.
-In a pre-registered, between-subject online experiment (N=759) and follow-up interviews (N=30), we compare (a) LLM-generated generic explanations, and (b) LLM-generated contextualized explanations.
-Our findings show that contextualized explanations (i.e., explanations that incorporate users’ past behaviors) effectively meet users’ cognitive needs while increasing users’ intentions to watch recommended movies.
-However, adding explanations offers limited benefits in meeting users’ utilitarian and affective needs, raising concerns about the proper design and implications of LLM-generated explanations.
-Qualitative insights from interviews reveal that referencing users’ past preferences enhances trust and understanding but can feel excessive if overused.
-Furthermore, users with more active and positive engagement with the recommender system and movie-watching get substantial gains from contextualized explanations.
-Overall, our research clarifies how LLM-generated recommendations influence users’ motivations and behaviors, providing valuable insights for the future development of user-centric recommender systems, a key element in social media platforms and online ecosystems.
+Large language models (LLMs) are increasingly integrated into recommender systems to generate natural language explanations for contextualized content.
+This study examines how different types of LLM-generated explanations shape user perceptions and behaviors in the context of movie recommendations.
+In a pre-registered, between-subject online experiment (N=759) and follow-up interviews (N=30), we compare LLM-generated (a) generic explanations, and (b) contextualized explanations.
+We find that contextualized explanations better address users’ cognitive and affective needs, and increase their intention to watch the recommended movies.
+Insights from interviews reveal that contextualization fosters trust but depends heavily on alignment with users’ genre and narrative preferences.
+Notably, including more historical movie references improves user perceptions and intent, though overly detailed explanations can feel redundant.
+Users with higher engagement or trust in the system benefit most from contextualization.
+These findings highlight the potential of LLMs to enhance user-centered recommendation experiences.
+We offer practical guidance for designing explanation strategies that balance personalization and cognitive effort to improve relevance, trust, and engagement across social media and entertainment platforms.
 Codehttps://anonymous.4open.science/r/Contextualizing-Recommendation-Explanations-With-LLMs-55D3
@@
 2022; Zhang and Chen 2020).
-The emergence of large language models (LLMs) presents potential for improving these recommendation explanations (Wu et al.
+The emergence of large language models (LLMs) presents potential for improving recommendation explanations (Wu et al.
 2024; Hou, Leach, and Huang 2024).
-Some studies have adopted LLMs to produce more natural and human-like explanations.
+Some studies have adopted LLMs to produce natural and human-like explanations.
 However, current efforts often rely on generic explanation templates (Li, Zhang, and Chen 2020).
@@
 Our goal is to examine how different LLM-generated recommendation explanation types resonate with specific user needs and, in turn, influence their consumption intentions.
-We design the following research question:
-How do different LLM-generated explanations respond to users’ diverse motivational needs (cognitive, affective, utilitarian) and affect their consumption intentions?
+Therefore, we designed the following research question:
+How do LLM-generated explanations respond to users’ diverse motivational needs (cognitive, affective, utilitarian) and affect their consumption intentions?
 Figure 1: Overview of the study process.
-In this study, we conducted a pre-registered, between-subject online experiment involving $N=759$ subjects from the U.S.
+In this study, we conducted a pre-registered, between-subject online experiment involving NN=759 subjects from the U.S.
 in the context of movie recommendations.
 [^1] The experimental protocol was pre-registered at OSF: https://osf.io/j46vx/?view_only=dc2b1f641b33407ab8168736a783055a By drawing upon the U&G framework, we analyze users’ perceptions and intentions towards three types of recommendations generated by GPT-4 Turbo: (1) recommendations with no explanations (NE), (2) recommendations with LLM-generated generic explanations (GE), and (3) recommendations with LLM-generated contextualized explanations (CE).
-We also drew upon user profile information and conducted follow-up interviews ($N=30$) to interpret our results more deeply.
+[^2] We focus on real-world recommendation tasks based on viewing histories without relying on explicit or private user preferences (e.g., as in cold-start problems of recommendation websites).
+As such, we use the term contextualized rather than personalized throughout the paper.
+This terminology aligns with related work (Bar, Maarouf, and Feuerriegel 2024) and reflects practical usage on platforms such as IMDb, Yelp, and Google Maps, where recommendations are often informed by user activity (e.g., viewing or browsing history) but not by direct feedback like ratings.
+We also drew upon user profile information and conducted follow-up interviews (NN=30) to interpret our results more deeply.
 To ensure the robustness of our study design and the generalizability of the results, we conducted supplementary experiments by using a different LLM (Claude 3.5 Sonnet) with an additional 140 users from the U.S.
@@
 2 Related Work
-2.1 Explanations in Recommender Systems
-In social media and online communities, recommendations help users to find items aligned with their preferences.
+2.1 LLM-Generated Explanations in RecSys
+In social media and online communities, recommendations help users find items aligned with their preferences.
+Initial research centered on improving predictive accuracy (Ma et al.
+2023; Wu et al.
+2023).
+However, growing attention to user trust and transparency has driven a shift toward explainable recommendations (Burkart and Huber 2021).
 Initial research centered on improving accuracy (Ma et al.
@@
 2023), but recent work emphasizes explainability to foster trust, transparency, and user satisfaction (Burkart and Huber 2021).
-Early approaches integrated explainability into model architectures, such as explainable matrix factorization leveraging user similarities (Abdollahi and Nasraoui 2017), tree-enhanced embeddings (Wang et al.
-2018), and attention-based neural models (Seo et al.
+Traditional recommendation techniques (e.g., matrix factorization with interpretable layers (Abdollahi and Nasraoui 2017) or attention-based neural models (Seo et al.
 2017; Chen et al.
-2018).
-More recent efforts have expanded explanation strategies, ranging from visual (Tsai and Brusilovsky 2019) and comparative methods (Le and Lauw 2021) to sentiment-oriented narratives (Chen, Yan, and Wang 2019), as well as traditional item-, content-, and popularity-based explanations (Sarwar et al.
-2001; Pazzani and Billsus 2007; Yang 2016).
-These approaches highlight the importance of making recommendations more understandable, interactive, and socially aligned to support user trust and transparency across online communities.
-Beyond these approaches, LLMs offer new opportunities for generating better explanations within the dynamic, user-driven landscapes (Mohanty 2024).
-By producing natural language explanations, LLMs can improve user satisfaction and trust, especially in domains where social context and user interaction are central (Chang, Harper, and Terveen 2016).
-For instance, LLMs have been used to create more detailed item descriptions (Acharya, Singh, and Onoe 2023) or to enable ChatGPT-based recommenders to generate user-friendly explanations (Deldjoo 2024).
-Recent research suggests that more contextualized, user-centered explanations could enhance trust and engagement (Silva et al.
-2024; Lyu et al.
-2024), reinforcing the value of LLM-driven strategies that align closely with the evolving social and informational needs of online communities.
-2.2 User Perceptions of Recommendation Explanations
+2018)) integrated explainability into model architectures.
+The advent of LLMs has made it easy to generate explanations in natural language more flexibly.
+Building upon template-based methods (Sarwar et al.
+2001; Pazzani and Billsus 2007; Yang 2016), studies have used LLMs to generate detailed item descriptions (Acharya, Singh, and Onoe 2023), visual explanations (Tsai and Brusilovsky 2019), sentiment-oriented narratives (Chen, Yan, and Wang 2019), and interactive user-friendly justifications (Deldjoo 2024).
+Previous works have also shown the unique potential of natural language explanations in mitigating information overload and facilitating community integration (Chang, Harper, and Terveen 2016).
+This growing trend of work underscores the potential for LLM-driven recommendation strategy in strengthening the recommendation processes.
+2.2 User Perceptions of Explanations
 Users increasingly rely on recommender systems to navigate overwhelming volumes of information.
-Research grounded in the U&G framework highlights that individuals engage with media to fulfill specific needs, which can be categorized as affective, cognitive, and utilitarian motivations (Katz, Blumler, and Gurevitch 1973; Korhan and Ersoy 2016).
-These motivations directly influence consumption intentions, as users are more likely to engage with media or systems that effectively address their various needs.
-High-quality explanations within recommender systems can significantly influence user engagement, satisfaction, and decision-making, particularly in socially dynamic environments where these systems satisfy informational and social needs (Mandl et al.
+Grounded in the Uses & Gratifications framework, individuals seek to satisfy their cognitive, affective, and utilitarian needs when engaging with social media content (Katz, Blumler, and Gurevitch 1973; Korhan and Ersoy 2016; Bashardoust, Feuerriegel, and Shrestha 2024).
+Accordingly, explanations that clarify why a recommendation is made can reduce cognitive effort, foster trust, and enhance decision confidence (Mandl et al.
 2011; Jameson et al.
 2015).
-For instance, explanations can enhance perceived usefulness and trust, aligning with motivation needs such as ease of use and task efficiency (Korhan and Ersoy 2016).
-Extant research also offers valuable insights into factors shaping user perceptions.
-For example, evidence shows that users may prefer human-generated recommendations, particularly for hedonic products (Wien and Peluso 2021).
-Similarly, explanations reflecting user feedback and personalization resonate with individuals who seek tailored, value-reinforcing content (Lu et al.
+Extant research also identifies several factors shaping user perceptions of explanations.
+For instance, users may prefer human-generated recommendations, particularly for hedonic products (Wien and Peluso 2021).
+Explanations that reflect prior user feedback or behavior tend to resonate with individuals seeking tailored, value-reinforcing content (Lu et al.
 2023).
-Differences in user profiles further emphasize the importance of detailed and specific explanations, particularly for users with high cognitive engagement who value transparency and trust (Chatti et al.
+Moreover, individual differences in user profiles, such as cognitive engagement level, highlight the importance of detailed and transparent justifications for users who prioritize control and understanding (Chatti et al.
 2022).
-Natural language explanations further enable systems to fulfill diverse needs, such as reducing cognitive effort, fostering trust, and enhancing community integration (Chang, Harper, and Terveen 2016).
-In parallel, researchers have developed evaluation metrics to address these user-centric demands beyond traditional measures like BLEU and ROUGE (Papineni et al.
-2002; Lin 2004).
-User-centered metrics now consider dimensions such as effectiveness, persuasiveness, satisfaction, and trust, emphasizing user needs in complex, community-driven contexts (Zhou et al.
-2021; Ehsan et al.
-2019; Hoffman et al.
-2023).
+Therefore, aligning explanation strategies with varied user motivations and perceptions is essential for designing effective and user-centric recommender systems.
+2.3 Empirical User Studies on Explanations
+Empirical research has analyzed how different explanations in recommender systems impact users (see Table 1).
+These studies typically use controlled experiments and surveys to evaluate diverse explanation strategies.
+Silva et al.
+(2024) conducted a user study (N=94) comparing ChatGPT-generated generic and personalized movie explanations.
+While both were well-received, personalization improved perceived effectiveness when recommendations were less relevant and enhanced persuasiveness in low-confidence scenarios.
+Similarly, Lubos et al.
+(2024) ran an online user study (N=97), finding that LLM-generated explanations outperformed template-based baselines in trust and satisfaction due to their contextual richness and fluency.
+In another user study of a fitness plan recommender system (N=341), Sun et al.
+(2023) showed that users trusted less of recommendations when explanations revealed reliance on their social-media friends’ activities, citing privacy and identity concerns.
+Instead, users preferred personalization based on their behavior.
+Together, these empirical findings provide insights into user perceptions and inform system design, but a comprehensive evaluation analyzing the cognitive, affective, and utilitarian needs is missing.
 2.4 Research Gap
@@
 3.1 Overview
-We conducted an online between-subject experiment, where subjects ($N=759$) were randomly assigned to one of three experimental conditions: (1) recommendations with no explanations (NE), (2) recommendations with LLM-generated generic explanations (GE), and (3) recommendations with LLM-generated contextualized explanations (CE).
+We conducted an online between-subjects experiment, where subjects (NN=759) were randomly assigned to one of three experimental conditions: (1) recommendations with no explanations (NE), (2) recommendations with LLM-generated generic explanations (GE), and (3) recommendations with LLM-generated contextualized explanations (CE).
 In the experiment, users first selected movies they had previously watched from a given set to establish the viewing history, based on which LLMs generated movie recommendations and explanations depending on the experimental condition.
@@
 These users still received remuneration after the exclusion.
-Eventually, due to our strict exclusion criteria, 759 of the 1,401 subjects qualified, exceeding our estimated sample size of 750 to ensure the results achieve the desired statistical power.
+Eventually, due to our strict exclusion criteria, 759 of the 1,401 subjects qualified, exceeding our estimated sample size of 750 to ensure the results achieved the desired statistical power.
 Among the 759 subjects, 387 are women, 370 are men, and two are non-binary.
@@
 We selected GPT-4 Turbo to generate movie recommendations for the main experiment.
-We conducted an additional experiment with Claude 3.5 Sonnet to validate the robustness of our study design and the generalizability of our findings (see supplementary materials).
-We precompiled all recommendations and explanations in a corpus to ensure that there was no variability in the outcomes across different LLM calls.
-To create the corpus, we first selected 10 movies with diverse topics from the IMDb Top 250 list [^4] https://www.imdb.com/chart/top/ to compile a movie shortlist $\mathcal{M}$.
-These movies were chosen based on their high ratings or popularity.
-We ensured that the shortlist had broad coverage of user viewing preferences.
-Subsequently, we implemented a sequential recommendation strategy following the algorithm in Table 5 in Appendix A.
-With this strategy, sujects with the same viewing history and assigned conditions received the identical recommendation..
+We conducted an additional experiment with Claude-3.5-Sonnet to validate the robustness of our study design and the generalizability of our findings (see Appendix C.1).
+We pre-compiled all recommendations and explanations in a corpus to ensure no variability in the outcomes across different LLM calls.
+The corpus is generated based on the pre-defined strategy and prompts below.
+Therefore, users with the same viewing history and assigned conditions received the identical recommendation explanation.
+Recommendation Strategy
+We implemented a sequential recommendation strategy following the algorithm in Table 5 in Appendix A.
+To reduce bias, users only rated movies they had not watched.
+For each user, we began the recommendation process with 10 distinct popular movies from the IMDb Top 250 list [^5] https://www.imdb.com/chart/top/.
+If the user had already watched a movie, we proceeded to new recommended movies until they reached an unwatched one.
+The recommended movies watched by users were also added to their viewing history and used to generate new recommendations.
+For example, if a user has watched Movie A before, LLMs will recommend Movie B.
+If the user has watched it as well, then LLMs will combine both movies in the explanation for Movie C.
+This process continued until either an unwatched recommendation was found or a maximum of 4 recommendations were generated.
+Rather than relying on explicit subjective feedback (e.g., likes or dislikes), we collected information about whether users had previously watched each movie.
+This design simulates a cold-start scenario (Lam et al.
+2008; Chaimalas et al.
+2023), where prior viewing history is unavailable and where the system starts with general popular choices before adapting to observed behavior.
+This approach mirrors real-world conditions, where explicit user feedback is often sparse or unavailable.
+Recommender websites (such as IMDb, Rotten Tomatoes, Google Maps, Yelp, and TripAdvisor) are typically user-independent and free of personalization, either because they have no access to user-specific information or personalization is unwanted.
+They only have access to consumption data from their registered users, that is, actual review contributors.
+For Yelp, it is estimated that only 0.67% of users are registered contributors, while the remaining users are visitors.
+[^6] See https://askwonder.com/research/percentage-yelp-users-mainly-write-reviews-restaurants-whrwmul7u, accessed March 12, 2025.
+Moreover, the main selling point of recommender websites like IMDb and Yelp is the comparison of products or services.
+As a result, users are not required to sign in to use the website.
+Forcing users to register on the website would probably be counterproductive and decrease activity, as users might not want to sign up due to privacy reasons.
+Consistent with these, we collect users’ viewing histories—rather than explicit preferences—to generate recommendations and explanations.
+Prompt Generation
 We followed previous practices (Giray 2023) when designing the prompts.
 Table 3 shows examples of contextualized and generic explanations.
-We can see a clear difference: while both explanations incorporate various elements (e.g., feature, popularity, similarity) to support the recommendations, the contextualized explanations naturally integrate the recommended movies with the user’s viewing history, explicitly highlighting how the user’s inputs shape the recommendation.
+All explanations were limited to 100 words, aligning with the length of typical movie descriptions.
+To manage reading speed and cognitive load, each explanation was designed to include around three key features, such as the director, cast, plot, or scenes.
+Further, the length of the explanations was similar across the different conditions to ensure comparability.
+The main difference between CE (contextualized explanations) and GE (generic explanations) is in the explicit connection: while both explanations incorporate various elements (e.g., feature, popularity, similarity) to support the recommendations, the contextualized explanations naturally integrate the recommended movies with the user’s viewing history, explicitly highlighting how the user’s inputs shape the recommendation.
 Meanwhile, the generic explanations focus exclusively on the recommended movies.
@@
 Step 11. Users receive a debrief of the experiment and are informed that LLMs power all movie recommendations.
-Afterward, we randomly reached out to $N=30$ users for the follow-up interviews.
+Afterward, we randomly reached out to NN=30 users for the follow-up interviews.
 [^8] Detailed interview questions in Appendix B.
@@
 We conducted analyses based on user perceptions in the survey.
-Each participant could provide up to 10 ratings, resulting in a total of 2,732 ratings from 759 participants in the main experiment (916 in NE, 898 in GE, and 918 in CE).
+Each participant could provide up to 10 ratings, resulting in a total of 2716 ratings from 759 participants in the main experiment (904 in NE, 911 in GE, and 901 in CE).
 We used hypothesis testing to identify differences in perceptions and consumption intentions across recommendation types, followed by mixed-linear regression to examine how user characteristics explain variation in both variables.
@@
 Since the data did not meet the normality assumption (Lilliefors 1967), we used a Mann–Whitney test with Holm correction to identify pairwise differences between conditions (McHugh 2011).
-Later, we use $\mu$ to report the mean.
+Later, we use $\mu$ to report the mean value.
 Mixed-effects linear regression.
@@
 Our final regression model is:
-$\emph{Y}_{ij}$ $=\beta_{0}+\beta_{1}\emph{Treatment}_{j}+\beta_{2}\emph{Age}_{j}+\beta_{3}\emph{Gender}_{j}+\beta_{4}\emph{Education}_{j}$
-$\quad+\beta_{5}\emph{MovieConsumption}_{j}+\beta_{6}\emph{RecSysFrequency}_{j}$ (1)
-$\quad+\beta_{7}\emph{RecSysUtility}_{j}+\beta_{8}\emph{DisplayOrder}_{j}+u_{0j}+\epsilon_{ij}$
-where ${Y}_{ij}$ denotes the dependent variables (i.e., Cognitive, Affective, Utilitarian, and UserIntent) for the $i$-th observation of the $j$-th subject, with intercepts $\beta_{0}$, subject-level random effects $u_{0j}$, coefficients $\beta_{1}$ through $\beta_{8}$, and an error term $\epsilon_{ij}$.
+$Y_{ij}$ $=\beta_{0}+\beta_{1}\emph{Treatment}_{j}+\beta_{2}\emph{Age}_{j}+\beta_{3}\emph{Gender}_{j}+\beta_{4}\emph{Education}_{j}$
+$\quad+\beta_{5}\emph{MovieConsumption}_{j}+\beta_{6}\emph{RecSysFrequency}_{j}$
+$\quad+\beta_{7}\emph{RecSysUtility}_{j}+\beta_{8}\emph{HistoryNum}_{j}+\beta_{9}\emph{DisplayOrder}_{j}$
+$\quad+\beta_{10}\emph{HistoryNum}_{j}\times\emph{MovieConsumption}_{j}$ (1)
+$\quad+\beta_{11}\emph{HistoryNum}_{j}\times\emph{RecSysUtility}_{j}+u_{0j}+\epsilon_{ij}$
+where ${Y}_{ij}$ denotes the dependent variables (i.e., Cognitive, Affective, Utilitarian, and UserIntent) for the $i$-th observation of the $j$-th subject, with intercepts $\beta_{0}$, subject-level random effects $u_{0j}$, coefficients $\beta_{1}$ through $\beta_{11}$, and an error term $\epsilon_{ij}$.
 3.6 Ethical Considerations
 We respect the privacy and agency of all subjects potentially impacted by this work and take steps to protect their privacy.
-The experiment design has obtained approval from the Ethics Commission at the University of Lausanne.
+The experiment design has obtained approval from the Ethics Commission at the University of (name anonymized for review).
 All data were collected anonymously, and any personally identifiable features were removed.
@@
 Below, we state the results of the hypothesis tests. The regression results are also shown in Table 4.
+Figure 2: Ratings by explanation type (shown variable: Cognitive, Affective, Utilitarian, and UserIntent). Significance levels between conditions are: $^{*}p<0.05$, $^{**}p<0.01$, $^{***}p<0.001$.
 4.1 Cognitive Needs
-We observed a significant difference (Holm-adjusted $p<0.05$) in the distribution of the aggregated Cognitive measure (i.e., Scrutability, Transparency, Trust) between NE ($\mu=4.87$, $95\%$ CI = [4.78, 4.95]) and CE ($\mu=5.22$, $95\%$ CI = [5.14, 5.30]).
-We also found a significant difference between GE and CE (Holm-adjusted $p<0.05$), with CE surpassing GE ($\mu=4.90$, $95\%$ CI = [4.81, 4.99]).
-However, there was no statistical difference between NE and GE (Holm-adjusted $p>0.05$), suggesting that generic explanations do not differ markedly from no explanations in meeting users’ cognitive needs.
-In the mixed-effects regression, both RecSysUtility ($\beta=0.319$, $p<0.05$) and MovieConsumption ($\beta=0.149$, $p<0.05$) demonstrated positive, significant effects on Cognitive, indicating that participants who perceive the recommender system as more useful and who watch movies more frequently tend to rate higher in Cognitive.
-In the follow-up interviews, several participants expressed increased trust and clarity when receiving the contextualized explanations.
-One participant noted: “When it told me the rationale for suggesting the movie, I could decide whether or not to watch it.” Another participant stated: “Including the movie I had watched and comparing the similarities made me trust the recommendation more.” Interestingly, some users also noted that references to movies they disliked helped them filter out unwanted content, improving their sense of control and overall transparency.
-These quantitative and qualitative findings suggest that contextualized explanations, by explicitly linking a user’s past and present preferences, can bolster trust and understanding in a way that generic or absent explanations do not.
-The more personal and concrete presentation of information serves users’ cognitive needs by explaining why some recommendations might be a good match, and others might not align with their tastes, ultimately leading to a deeper sense of scrutiny, transparency, and trust.
-Figure 2: Cognitive ratings by explanation type (shown variable: Cognitive). Significance levels between conditions are indicated: ${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$.
+We observed a significant difference (Holm-adjusted $p<0.05$) in the distribution of the aggregated Cognitive measure between NE ($\mu$=4.94), GE ($\mu$=4.97), and CE ($\mu$=5.38).
+Pairwise tests showed that CE significantly outperformed both NE and GE (Holm-adjusted pp<0.05), indicating that contextual explanations better support users’ cognitive understanding.
+However, there was no statistical difference between NE and GE (Holm-adjusted pp>0.05), suggesting that generic explanations offer no additional cognitive benefit over providing no explanation.
+In the mixed-effects regression, both RecSysUtility ($\beta$=0.304, pp<0.05) and MovieConsumption ($\beta$=0.150, pp<0.05) demonstrated positive, significant effects on Cognitive, indicating that participants who perceive the recommender system as more useful and who watch movies more frequently tend to rate higher in Cognitive.
+Besides, more movie contexts ($\beta$=0.048, pp<0.05) in the explanation also better satisfy users’ cognitive needs.
+In the interviews, several participants clearly expressed their trust and commended the clarity when receiving contextualized explanations.
+“The thing that makes me trust the explanation was reading how the description had brought up other titles of movies that I saw in the recommendations, that had true things about the characters, and how the movies could be similar or how I could watch them already.” (Female, 29)
+“They made the details of the movies sound more exciting and interesting and gave you more of a full rundown of theories so you don’t think it’s going to be one thing and turn out another.” (Female, 31)
+On the other hand, sometimes users disliked the recommendations, but they still helped to filter out unwanted content, improving their sense of control and overall transparency.
+“The movies it recommended didn’t interest me, so I don’t think I’ll be watching them. Felt like I understood the movies weren’t for me.” (Male, 32)
+“I trusted the recommended movie, but I made my own mind up after reading the explanation.” (Male, 44)
 4.2 Affective Needs
-We found no overall difference in the aggregated Affective measure (i.e., Persuasiveness, Satisfaction) among NE ($\mu=4.97$, $95\%$ CI = [4.88, 5.05]), GE ($\mu=4.92$, $95\%$ CI = [4.83, 5.01]), and CE ($\mu=5.02$, $95\%$ CI = [4.93, 5.11]).
-However, examining each subdimension showed that Persuasiveness differed significantly between GE and CE (Holm-adjusted $p<0.05$), indicating that contextualized explanations can make recommendations more compelling than generic explanations, even if this advantage does not fully carry over to overall perceived affective needs.
-The mixed-effects regression highlights two user-engagement variables, RecSysUtility ($\beta=0.311$, $p<0.05$) and MovieConsumption ($\beta=0.167$, $p<0.05$), as positive, significant predictors of affective responses.
-This suggests that users who view the recommender system as helpful and frequently watch movies are more likely to feel emotionally drawn in, regardless of the explanation style.
-Our interviews offer deeper insights into how these affective needs take shape.
-Several participants emphasized that a good explanation heightened their eagerness to watch the movie: “They made me want to watch them more than I would have before with their explanations.” Others found “overly pushy” or “over the top” descriptions off-putting, highlighting the delicate balance between building excitement and overselling the film.
-Some appreciated specific references to actors or plot details, while others disliked spoiler-like content, favoring a brief, suggestive teaser.
-These varying reactions prove that users’ emotional engagement hinges on contextual cues and personal taste, mood, and openness to promotional language.
-These findings imply that satisfying affective needs, such as feeling entertained, inspired, or emotionally “in tune”, goes beyond simply adding contextual detail.
-Users’ enjoyment levels are tightly bound to their existing trust in the system, general interest in watching movies, and thresholds for promotional or spoiler content.
-Consequently, while contextualized explanations enhance persuasiveness for some users, effectively catering to affective motivations requires a more nuanced approach that recognizes diverse emotional triggers and preferences.
-Figure 3: Affective ratings by explanation type (shown variable: Affective).
+For the Affective dimension, CE ($\mu$=5.19) was rated significantly higher than both NE ($\mu$=5.05) and GE ($\mu$=4.99) (Holm-adjusted pp<0.05).
+Still, no significant difference emerged between NE and GE (Holm-adjusted pp>0.05).
+The mixed-effects regression highlights two user-engagement variables, RecSysUtility ($\beta$=0.295, pp<0.05) and MovieConsumption ($\beta$=0.175, pp<0.05), as positive, significant predictors of affective responses.
+This suggests that, regardless of the explanation style, users who view the recommender system as helpful and frequently watch movies are more likely to have positive feedback in affective needs.
+Besides, more movie context ($\beta$=0.069, pp<0.05) in the explanation has a positive effect as well.
+From the interview, we observed that the movie and explanation quality are important to users’ affection, no matter whether it contains the context.
+This also explains why users’ satisfaction is fairly similar when receiving CE or GE.
+Regarding the textual quality, the output of LLMs is consistent regardless of the explanation types.
+“Because I am a movie buff, and frankly, the descriptions were very well laid out so that it gave me enough information about the plot but also left enough unsaid that now I have to find out.” (Male, 39)
+“Some of the word descriptions were a little over the top. The description was a little graphic.” (Female, 68)
+“The explanation was well written, so it gave away a little more than I would normally like, but was just as good as a trailer in my mind.” (Male, 39)
+Figure 3: Radar chart of average user perception ratings in different dimensions by explanation types.
 4.3 Utilitarian Needs
-There is no overall difference in the aggregated Utilitarian measure (i.e., Effectiveness, Efficiency, Helpfulness) across NE ($\mu=5.24$, $95\%$ CI = [5.16, 5.32]), GE ($\mu=5.27$, $95\%$ CI = [5.19, 5.36]), and CE ($\mu=5.27$, $95\%$ CI = [5.18, 5.35]).
-However, a closer examination at Helpfulness revealed that NE ($\mu=5.07$, $95\%$ CI = [4.98, 5.17]) versus CE ($\mu=5.23$, $95\%$ CI = [5.13, 5.33]) was significant (Holm-adjusted $p<0.05$), suggesting that contextualized explanations can offer an advantage over having no explanation in practice by helping users assess whether a movie meets their needs.
-By contrast, there was no difference between GE and CE in Helpfulness, indicating that once any form of explanation is provided, adding contextual details may not further enhance users’ perceived practical value in selecting movies.
-The mixed-effects regression showed RecSysUtility ($\beta=0.292,p<0.05$) and MovieConsumption ($\beta=0.181,p<0.05$) both positively influenced Utilitarian, again implying that users who closely engage with recommender systems and movies, tend to see greater practical value in the recommendations regardless of explanation style.
-Female participants also reported slightly higher utilitarian scores (versus male, $p<0.05$).
-Our interview provides further evidence.
-Several participants appreciated having clear, concise details about the plot or genre, emphasizing how these insights save time in deciding whether a movie fits their tastes: “The more it detailed the movie, the more it helped me decide.” Some also suggested that referencing both liked and disliked movies aided them in filtering out poor matches and reinforcing good ones: “I was able to assess whether I wanted to watch it, based on what I liked before, and also on what I knew I didn’t enjoy.” Others felt the explanations were too vague or enthusiastic, hindering their ability to gauge true quality.
-Although contextualized explanations may slightly outperform no explanations in offering helpful guidance, overall utilitarian gains do not appear large when considering Effectiveness or Efficiency.
-These findings suggest that users’ utilitarian needs stem less from how the explanation is framed and more from whether the system is already perceived as useful and aligns with existing consumption habits.
-Nonetheless, highlighting specific points of similarity or dissimilarity can provide an additional benefit in helping users quickly identify which recommendations are worth pursuing and which to skip.
-Figure 4: Utilitarian ratings by explanation type (shown variable: Utilitarian).
+In the Utilitarian dimension, CE ($\mu$=5.44) was rated significantly higher than NE ($\mu$=5.32) (Holm-adjusted pp<0.05).
+There was no significant difference between NE and GE or CE and GE ($\mu$=5.34).
+The mixed-effects regression showed RecSysUtility ($\beta$=0.271, pp<0.05) and MovieConsumption ($\beta$=0.172, pp<0.05) both positively influenced Utilitarian, implying that users who closely engage with recommender systems and movies tend to see greater practical value in the recommendations regardless of explanation style.
+Female participants also reported higher utilitarian scores (versus male, $\beta$=0.292, pp<0.05).
+More movie context increases utilitarian value ($\beta$=0.044, pp<0.05).
+In general, users felt that these explanations save time in deciding whether a movie fits their tastes.
+However, we observed that some users appreciated more details to support their decisions, while some preferred the explanations to be simplified.
+“If I couldn’t clearly understand the explanation of the film, I thought that I was most likely not going to enjoy the movie.
+I trust my first instincts on things because it has saved me a lot of time in the past, by allowing me to avoid seeing a movie that would only waste time in my too short life.” (Female, 37)
+“Sometimes too much is overwhelming.” (Male, 37)
+“The more detailed movie, the more it helped me decide.”(87, Female)
 4.4 Consumption Intentions
-Knowing users’ consumption intentions (measured by UserIntent) helps us understand how explanations translate into actual viewing decisions.
-Pairwise comparisons confirm that CE ($\mu=4.70$, $95\%$ CI = [4.59, 4.81]) significantly differs from GE ($\mu=4.31$, $95\%$ CI = [4.18, 4.43]) (Holm-adjusted $p<0.05$).
-The contrasts between NE ($\mu=4.53$, $95\%$ CI = [4.42, 4.65]) and CE, as well as NE and GE, are also significant (Holm-adjusted $p<0.05$).
-This pattern suggests that the existence of explanations and contextualization fosters greater motivation to watch a recommended movie.
-Mixed-effects regression indicates that CE increases consumption intentions by 0.240 points on average ($\beta=0.240,p<0.05$) compared to GE.
-RecSysUtility ($\beta=0.336,p<0.05$) and MovieConsumption ($\beta=0.143,p<0.05$) both increase the likelihood of watching the recommended title.
-Meanwhile, older participants ($\beta=-0.139,p<0.05$) and female users ($\beta=-0.227,p<0.05$) expressed slightly lower consumption intentions.
-The qualitative interview responses underscore how alignment with personal taste can motivate users to watch the recommended movies.
-Several participants mentioned that recognizable themes, plot twists, or well-known actors greatly influenced their choices.
-One participant explained: “An interesting twist or variation from an expected plot development intrigues me, as does the presence of an actor I like.” Meanwhile, others highlighted that the mention of disliked or irrelevant elements quickly led them to dismiss a recommendation: “Some movies I simply will not watch … if it’s not my kind of genre, no explanation can convince me.” In some instances, participants weighed social factors such as the opinions of friends and family members, underscoring how shared viewing decisions can override individual preferences if, for example, a partner or group prefers a different type of film.
-These results imply that while contextualized explanations provide an edge over generic ones, individual beliefs about system helpfulness and prior consumption habits remain critical in determining whether users actually watch a recommended movie.
-Explaining “why this movie” helps, but so does the user’s fundamental perception of the recommender’s reliability and personal engagement with the medium.
-Ultimately, consumption intentions emerge from a synthesis of explanation quality, personal taste, and social or experiential factors, illustrating that users weigh both the persuasive power of contextualization and their unique circumstances.
-Figure 5: Consumption intentions by explanation type (shown variable: UserIntent). Significance levels between conditions are indicated: ${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$.
+We found significant differences in the consumption intention across all conditions.
+CE ($\mu$=4.85) outperformed GE ($\mu$=4.36) and NE ($\mu$=4.61) (Holm-adjusted pp<0.05).
+GE was also rated significantly lower than NE (Holm-adjusted pp<0.05).
+Mixed-effects regression indicates that RecSysUtility ($\beta$=0.292, pp<0.05) and MovieConsumption ($\beta$=0.125, pp<0.05) both increase the likelihood of watching the recommended title.
+Older participants ($\beta$=-0.147, pp<0.05) expressed lower consumption intentions.
+Besides, more movie context ($\beta$=0.101, pp<0.05) in the explanation has a positive effect as well.
+Though users have overall positive feelings, their consumption intent is more attached to their taste.
+Many interviewees explicitly expressed that, if the movie is not their type, they will not watch it.
+This is not a consequence of our findings but of the accuracy of the algorithm behind the recommender systems.
+Hence, this still confirms the value of contextualization in motivating users’ decisions.
+“The factor that motivated me the most in deciding whether or not to watch the movie was the correlation factor between the new movie and movies I have watched in the past.”(Female, 20)
+“Sometimes the explanation was so intriguing I felt as though I wanted to know more about the movie.
+If the explanation is so boring, I don’t finish it, I’m most likely not going to watch the movie.” (Female, 37)
+“If it seemed like the movie was for children, it was a no for me.”(Male, 32)
+“This is becoming a little redundant, but again it explained it to the point I could see in my head, the picture and the action happening.
+Plus, anything with Will Smith has got to be awesome.” (Female, 61)
 Cognitive Affective Utilitarian UserIntent
 Intercept
-4.870***
-4.906***
-5.084***
-4.642***
+4.959***
+5.030***
+5.172***
+4.849***
 Treatment (GE)
-0.030
 0.010
-0.082
---0.180
+-0.032
+0.064
+-0.232
 Treatment (CE)
-0.366***
-0.125
-0.118
-0.240*
+0.418***
+0.166
+0.165
+0.277*
 Gender (Female)
-0.094
-0.134
-0.284***
---0.227*
+0.120
+0.160
+0.292***
+-0.181
 Age
---0.061
---0.076
-0.025
---0.139*
+-0.075
+-0.089*
+0.013
+-0.147*
 Education
-0.049
+0.038
+0.028
+0.003
+0.040
+RecSysUtility
+0.304***
+0.295***
+0.271***
+0.292***
+RecSysFrequency
+0.019
+-0.004
+-0.023
+0.132
+MovieConsumption
+0.150**
+0.175***
+0.172***
+0.125*
+DisplayOrder
+0.040
+0.035
+0.017
+0.115**
+HistoryNum
+0.048**
+0.069***
+0.044**
+0.101***
+HistoryNum $\times$ MovieConsumption
+-0.028
+0.007
+-0.032*
 -0.047
+HistoryNum $\times$ RecSysUtility
 -0.019
-0.060
-RecSysUtility
-0.319***
-0.311***
-0.292***
-0.336***
-RecSysFrequency
+-0.036*
 -0.013
---0.004
---0.043
-0.106
-MovieConsumption
-0.149***
-0.167***
-0.181***
-0.143*
-DisplaypOrder
---0.018
---0.046
---0.031
---0.022
-Significance levels: ${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$
-Table 4: Mixed-effect linear regression results. Coefficients are shown with standard errors in parentheses.
-5 Supplementary Experiments
+-0.054
+Table 4: Mixed-effects linear regression results. Coefficients are reported with standard errors in parentheses. Significance levels: $^{*}p<0.05$, $^{**}p<0.01$, $^{***}p<0.001$.
+4.5 Interpretations
+Based on the results above, we observed an overall advantage of LLM-generated contextualized explanations in addressing users’ motivational needs and increasing their consumption intentions.
+Figures 2 and 3 show that contextualized explanations successfully address users’ diverse needs.
+However, a key limitation remains: the explanations do not always align with users’ literary and genre preferences.
+Although we asked participants to evaluate the explanations rather than the recommended movies, interviews revealed that many users still evaluated the two together, resulting in lower scores for persuasiveness and consumption intention.
+When users turn to recommender systems, they often have specific goals, such as finding a movie to watch or exploring new options.
+Once the system delivers relevant recommendations, users may feel that their immediate needs are already satisfied.
+This likely explains why participants rated effectiveness, efficiency, and satisfaction high, even in the absence of explanations.
+Together, our findings highlight the value of LLM-generated explanations in building transparency and trust.
+However, to truly enhance user experience, explanations need to resonate with users’ tastes and expectations, especially in subjective domains like entertainment.
+Demographics and prior engagement shape user perception: Based on the results in Table 4, users’ prior engagement significantly shapes their evaluation of explanations.
+Participants who consumed more movies rated explanations more positively across all dimensions, suggesting that domain familiarity enhances appreciation of contextualized content.
+Similarly, higher perceived usefulness of recommender systems strongly gave favorable evaluations.
+Demographic effects were more selective.
+Female users gave higher utilitarian ratings, while older users rated affective and intention outcomes slightly lower.
+These findings suggest that both system engagement and user characteristics should be considered when designing a user-friendly system.
+More viewing history contexts, better perceptions: Table 4 reveals that the number of viewing histroy (HistoryNum) positively influenced users’ evaluations in perception and intention.
+As detailed in Section 3.3, we limited the number of contexts maximum of four, which was guided by typical movie description lengths, native reading speeds, and cognitive load thresholds.
+Importantly, the non-significant interaction effects between HistoryNum and both MovieConsumption and RecSysUtility suggest that the positive impact of including more movie contexts in explanations holds consistently, regardless of users’ movie consumption frequency or perceptions of the recommender system’s helpfulness.
+This implies that richer contextualization can enhance user evaluations without being contingent on prior engagement.
+Although richer context tends to improve perception and intent, the effects, based on the coefficients in the regression results, remain relatively small.
+Some users also flagged certain explanations as redundant or overly detailed.
+To prevent cognitive overload, we should carefully calibrate how much context to include, ensuring that explanations remain concise yet informative.
+5 Heterogeneity Analysis
+5.1 Claude 3.5 Sonnet
 To supplement the main experiment, we expanded the experiment to a larger scope by leveraging another LLM (Claude 3.5 Sonnet).
@@
 We generated another recommendation corpus using Claude 3.5 Sonnet and replicated the experiments on the U.S.
-samples ($N=140$, 71 are women, 68 are men, and one is non-binary).
+samples (NN=140, 71 are women, 68 are men, and one is non-binary).
 The results are aligned with our main findings, validating the advantages of LLMs in generating contextualized explanations in terms of meeting users’ motivational needs and increasing their consumption intent.
-The results and interpretations are provided in Appendix C.
+The results and interpretations are provided in Appendix C.1.
+5.2 Movie Genre
+To further understand how user perceptions vary across recommendation contexts, we examined the role of movie genre, specifically comedy and adventure, with supplementary mixed-effects regression models (see Equation C.2 in Appendix C.2).
+As shown in Table 7, comedy movies were consistently rated lower than adventure movies across affective, utilitarian, and consumption intention outcomes.
+We also observed that these genre effects were not moderated by explanation types.
+These supplementary results highlight genre as a key factor influencing user responses to LLM-generated recommendations.
+The influence of movie genre remains an important area for further investigation.
 In summary, the supplementary experiment yielded results that supported the main experiment’s findings and conclusions, thereby validating the robustness of our study design and the generalizability of the findings.
@@
 6.1 Implications of Findings
-Recommender systems play a central role in shaping user decisions, particularly influential in the context of web and social media.
-Given the far-reaching implications of these systems, there is growing demand from users and regulators for transparency, emotional connection, or efficiency in decision-making.
+Recommender systems play an important role in shaping user decisions, particularly influential in the context of web and social media.
+Given the far-reaching implications of these systems, there is growing demand from users and regulators for transparency, emotional connection, and efficiency in decision-making.
 This highlights the value of high-quality and context-aware recommendation explanations.
@@
 Our study demonstrates that LLM-generated contextualized explanations can overcome these limitations by satisfying users’ cognitive needs, ultimately increasing their consumption intentions.
-Also, our findings reveal the insufficiency of LLM-generated explanations in meeting users’ utilitarian and affective needs, which raises concerns about the proper design and implications of LLM-generated personalized content.
-Benefits of contextualization.
-Our results show that contextualized explanations significantly satisfy users’ needs and their intention to watch the recommended movies.
-This is particularly evident in users’ cognitive needs (Scrutability, Transparency, Trust), where contextualization not only promotes a clearer understanding of why specific titles are recommended but also encourages users to scrutinize recommendations more carefully.
-The increase in trust and perceived transparency shows the importance of tailoring explanations to individual needs.
-Offering personally relevant and meaningful recommendations can encourage users to act on these suggestions, especially in entertainment and e-commerce settings, where a personalized touch can be key to customer retention.
-LLMs in the recommendation.
-As LLMs evolve, their capacity to generate richer, more personalized explanations will continue to grow (Feuerriegel et al.
-2025).
-This capability is particularly relevant for companies like Amazon and Netflix, where personalization is a competitive differentiator (Burroughs 2019).
-Our findings highlight how integrating LLMs into recommender systems can satisfy users’ cognitive needs and increase their consumption intention.
-Beyond driving engagement, context-aware explanations may help companies fulfill emerging regulatory requirements, such as the EU’s AI Act, which calls for transparency and the “right to explanation” in AI-driven decisions (Veale and Borgesius 2021; You and Suh 2024).
-System utility and positive feedback loops.
-Our study suggests that a user’s active and positive engagement with the recommender system amplifies the benefits of contextualization.
-When participants view the system as helpful and relevant to their interests, context-aware explanations further boost their ratings of recommendation quality, trust, and helpfulness.
-This synergy can create a positive feedback loop: as users find the system increasingly beneficial, they engage more deeply, and contextualized explanations, in turn, reinforce these favorable perceptions.
-Moreover, frequent movie watchers also exhibit higher cognitive, affective, and utilitarian evaluations.
-Increasing user engagement through personalized interfaces, easy-to-use features, or curated content could, therefore, magnify the impact of contextualized explanations (Lacker and Way 2024).
-Companies seeking to cultivate sustained user relationships can harness this dynamic by combining context-rich recommendation details with user education or incentives that further encourage exploration and use of the system.
-Avoiding redundant information.
-An interesting insight from the experiment is that some participants found overly “pushy” or “over the top” descriptions off-putting.
-In contrast, others appreciated detailed explanations.
-Although contextualized explanations can meet various user needs, avoiding irrelevant or excessive information that increases cognitive load is critical.
-Our findings suggest that, for specific users, no explanation is preferable to offering a generic one.
-This reflects a balancing act: well-crafted, context-specific explanations can significantly improve user perceptions, but lengthy or poorly targeted explanations can become a distraction.
-In industries such as travel and finance, where decisions can be complex, selective use of contextualized explanations could help users make informed choices without overwhelming them.
-By highlighting only the most pertinent details, companies can streamline the decision-making process, thus promoting higher engagement and satisfaction.
+Also, our findings reveal the insufficiency of LLM-generated explanations in meeting users’ utilitarian and affective needs, which raises concerns about the proper design and implications of LLM-generated contextualized content.
+Build trust, but risk misrepresenting preferences: Our results show that contextualized explanations significantly address users’ cognitive needs, particularly in terms of Transparency, Trust, and Scrutability.
+By referencing users’ prior viewing history, LLM-generated explanations appear more tailored and meaningful, encouraging deeper engagement with recommendations.
+Nevertheless, we also observed that LLMs tend to interpret viewing histories as an implicit preference signal.
+While this can increase perceived personalization, it also introduces a risk of over-interpretation, which is reflected in user interviews.
+In practice, real-world recommender systems often infer preferences from behavioral signals (e.g., clicks, watch time) rather than explicit feedback, which may not always align with users’ true interests or sentiments.
+LLMs, when generating explanations based on such data, may reinforce these weak or ambiguous signals, thereby creating a disconnect between the explanation and the user’s actual intent.
+This misalignment risks undermining the trust that contextualized explanations build.
+When users feel the system attributes preferences they do not hold, trust may erode over time.
+Therefore, in real-world deployments, systems that leverage LLMs for explanation generation should account for this interpretive bias and consider more nuanced strategies for referencing user history, particularly when preferences are inferred rather than explicitly stated.
+Opportunities and challenges for the movie industry and users: Our findings show that LLM-generated movie recommendations significantly increase users’ intention to engage with suggested content.
+Effective recommendations can be generated by leveraging typical user viewing histories and state-of-the-art LLMs—without requiring additional user input or behavioral interventions.
+Such ease of implementation, coupled with LLMs’ ability to produce highly contextualized content, makes them valuable tools for streaming services and other media platforms.
+Integrating LLMs into recommendation systems offers the potential to enhance operational efficiency but introduces significant costs that vary by deployment strategy.
+Using commercial APIs such as OpenAI’s GPT-4o mini, priced at \$0.3 per million input tokens and \$1.2 per million output tokens [^10] OpenAI API Pricing: https://openai.com/api/pricing/, a typical user interaction (200 input and 100 output tokens) costs approximately \$0.00018.
+If each of Netflix’s 301.6 million global subscribers received 50 recommendations daily, the resulting API cost would amount to approximately \$2.7 million per day.
+These costs must be weighed against potential revenue gains.
+In subscription-based models like Netflix, the average revenue per user in North America was about \$17.17 in 2024 (Singh 2025).
+In advertisement-driven platforms like YouTube, creators typically earn \$5 to \$7 per 1,000 ad views (Marshall 2024).
+Even modest improvements in user engagement could yield substantial returns, emphasizing the importance of careful cost-benefit analysis when scaling LLM integration.
+Further, while LLMs present substantial opportunities to enhance recommendation quality and operational efficiency, successful deployment requires careful management of associated technical, economic, and ethical challenges.
+In particular, LLM-generated recommendations carry risks related to biased outputs (Feuerriegel et al.
+2023; Bashardoust et al.
+2024).
+In movie recommendation contexts, biases may manifest as an underrepresentation of specific genres, cultures, or demographic perspectives.
+For example, the corpus generated in our study tended to favor Western-centric narratives disproportionately.
+Additionally, perceived biases or cultural insensitivities arising from skewed training data or adversarial prompt manipulation can erode user trust and pose significant reputational risks to platforms.
 6.2 Limitations
 While our study offers valuable insights into user perceptions of LLM-generated explanations for movie recommendations, several limitations should be noted.
-First, we primarily recruited English-speaking users. Future research should include more diverse populations to strengthen the generalizability of our findings.
-Second, the effectiveness of the explanations may have been influenced by the performance of the underlying LLMs.
-Although we used state-of-the-art models (GPT-4 Turbo and Claude 3.5 Sonnet), we should notice that future advancements in LLMs could further shape or change user perceptions.
-Third, our distinction between contextualized and generic explanations may oversimplify the explanation types.
-Future studies should explore a broader variety of explanation strategies to better understand their impact in different contexts and on diverse user groups.
-Lastly, our experimental design did not fully capture the dynamic real-world user interactions with recommender systems.
-User perceptions might change after consuming the recommended movies, which is not covered by this study.
-This highlights the necessity of conducting the user study in real scenarios.
+First, our findings may be limited in generalizability due to both the participant sample and the underlying models.
+We primarily recruited English-speaking users, which may not capture the full range of cultural and linguistic differences in user perceptions.
+Additionally, the effectiveness of the explanations likely depends on the capabilities of the underlying LLMs.
+Although we used state-of-the-art models (GPT-4 Turbo and Claude 3.5 Sonnet), future advancements may further influence how users interpret and respond to recommendations and explanations.
+Second, our study may oversimplify the explanation generation process.
+Future research should explore various explanation strategies to better understand their effectiveness across different contexts.
+While we employed a typical cold-start recommendation scenario, providing additional user information, such as demographics or expressed preferences, can more efficiently reveal LLMs’ strong contextualization capabilities.
+Lastly, our experimental design does not account for the longitudinal dynamics of real-world user interactions with recommender systems.
+While we capture users’ immediate perceptions of explanations and recommendations, these impressions may evolve as users continue to engage with the recommended content over time.
+Additionally, due to our control over user inputs, participants did not have the opportunity to express specific or niche movie preferences to the LLMs, which may play a significant role in shaping their perceptions.
+Moreover, the broader long-term effects, such as the cumulative impact on content consumption patterns and industry outcomes, remain open for future investigation.
 Despite these limitations, our study provides a strong foundation for further research into the role of LLM-generated explanations in recommender systems, emphasizing the importance of contextualization in improving user experience and engagement.
@@
 Our findings show how and why recommendation explanations meet cognitive, affective, and utilitarian motivations under the U&G perspective.
-Our study confirms that LLM-generated contextualized explanations effectively address users’ cognitive needs and increase their consumption intentions.
+Our study confirms that LLM-generated contextualized explanations effectively address users’ cognitive and affective needs and increase their consumption intentions.
 Notably, recommender systems should also avoid overwhelming users with irrelevant information that can hinder decision-making or erode trust.
@@
 3     if $k=1$ then
-4      Generate $r_{i1}$ from $m_{i}$
-     $r_{ik}$: The $k$-th recommendation for the $i$-th movie
+4      Generate $r_{i,1}$ from $m_{i}$
+     $r_{i,k}$: The $k$-th recommendation for the $i$-th movie
 5     else
-6      Generate $r_{ik}$ from $\{m_{i},r_{i1},\dots,r_{i(k-1)}\}$
-     Previous recommendations are incorporated to generate $r_{ik}$
+6      Generate $r_{i,k}$ from $\{m_{i},r_{i,1},\dots,r_{i,(k-1)}\}$
+     Previous recommendations are incorporated to generate $r_{i,k}$
 7     end if
-8     Generate $e_{ikc}$ and $e_{ikg}$ for $r_{ik}$
-    $e_{ikc}$: Contextualized Explanation for $r_{ik}$
-    $e_{ikg}$: Generic Explanation for $r_{ik}$
-9     Save $r_{ik}$, $e_{ikc}$, and $e_{ikg}$ in $\mathcal{C}$
+8     Generate $e_{i,k,c}$ and $e_{i,k,g}$ for $r_{i,k}$
+    $e_{i,k,c}$: Contextualized Explanation for $r_{i,k}$
+    $e_{i,k,g}$: Generic Explanation for $r_{i,k}$
+9     Save $r_{i,k}$, $e_{i,k,c}$, and $e_{i,k,g}$ in $\mathcal{C}$
     Corpus $\mathcal{C}$ stores the recommendations and their explanations
@@
 Table 5: Recommendation corpus preparation steps.
-The corpus $\mathcal{C}$ contains recommendations $r_{ik}$ for each movie $m_{i}\in\mathcal{M}$, with contextualized explanations ($e_{ikc}$) and generic explanations ($e_{ikg}$) for each recommendation.
+The corpus $\mathcal{C}$ contains recommendations $r_{i,k}$ for each movie $m_{i}\in\mathcal{M}$, with contextualized explanations ($e_{i,k,c}$) and generic explanations ($e_{i,k,g}$) for each recommendation.
+Table 5 details the procedure we use to construct the recommendation corpus $\mathcal{C}$ from a given movie shortlist $\mathcal{M}$.
+For each movie $m_{i}\in\mathcal{M}$, the system generates a sequence of four recommendations, indexed by $k=1$ to $4$.
+At each step:
+  • The variable $m_{i}$ denotes a single movie from the input shortlist.
+  • The index $k$ indicates the recommendation order (which, equivalently, is the number of movies in the users viewing history used to produce the recommendation).
+  • The system generates $r_{i,k}$, the $k$-th recommendation for movie $m_{i}$.
+For the first recommendation ($k=1$), the system conditions only on the original movie $m_{i}$.
+For subsequent recommendations ($k>1$), it incorporates $m_{i}$ and all previously generated recommendations ($r_{i,1},\dots,r_{i,(k-1)}$) to produce $r_{i,k}$.
+This iterative approach models a realistic, context-aware recommendation process.
+After generating each recommendation $r_{i,k}$, the system produces two types of explanations for users in different conditions:
+  • $e_{i,k,c}$: a contextualized explanation that references the earlier items in the recommendation sequence;
+  • $e_{i,k,g}$: a generic explanation that relates $r_{i,k}$ to the original movie $m_{i}$ without incorporating context.
+Each recommendation $r_{i,k}$ and its corresponding explanations ($e_{i,k,c}$ and $e_{i,k,g}$) are added to the final corpus $\mathcal{C}$.
+This procedure ensures that the corpus captures both sequential recommendation dynamics and consistency across explanation styles.
 Appendix B Interview Questions
@@
   • “How do you think your profile, such as your demographics or viewing habits, influenced how you felt about the explanation you received?”
-Appendix C Supplementary Experiment (with Claude 3.5 Sonnet)
+Appendix C Supplementary Experiments
+C.1 Clade 3.5 Sonnet
 We replicated the experiment on a smaller U.S.
-sample ($N=140$) using Claude 3.5 Sonnet.
+sample (NN=140) using Claude 3.5 Sonnet.
 The results align with our main study, showing that contextualization improves overall explanation quality, particularly in terms of persuasiveness, transparency, and trust.
-Moreover, contextualized explanations increase users’ consumption intentions.
-Additionally, the positive effect of RecSysUtility is maintained, even with another underlying LLM instead of GPT-4 Turbo.
+Besides, contextualized explanations increase users’ consumption intentions.
 While some minor discrepancies emerged between this supplementary experiment and the main experiment, the core findings remain consistent.
-This supports the generalizability of our study’s conclusions.
-Figure 6, Figure 7, Figure 8, and Figure 9 show the user ratings of motivational needs and consumption intentions in this experiment.
-Figure 6: Cognitive ratings by explanation type (shown variable: Cognitive). Significance levels between conditions are indicated: ${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$.
-Figure 7: Affective ratings by explanation type (shown variable: Affective).
-Figure 8: Utilitarian ratings by explanation type (shown variable: Utilitarian).
-Figure 9: Consumption intentions by explanation type (shown variable: UserIntent). Significance levels between conditions are indicated: ${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$.
+Due to the limited sample size, some of the results are insignificant.
+Overall, the results of the supplement experiment support the generalizability of our study’s conclusions.
+Figure 5 shows the user ratings of motivational needs and consumption intentions.
+Figure 4 presents a more detailed comparison across dimensions.
+Table 6 shows the regression results.
+Figure 4: Radar chart of average user perception ratings in different dimensions by explanation types (Model: Claude 3.5 Sonnet).
+Figure 5: Ratings by explanation type (shown variable: Cognitive, Affective, Utilitarian, and UserIntent).
+Significance levels between conditions are indicated: $^{*}p<0.05$, $^{**}p<0.01$, $^{***}p<0.001$.
 Cognitive Affective Utilitarian UserIntent
 Intercept
-4.823***
-4.807***
-5.103***
-4.559***
+4.955***
+4.935***
+5.280***
+4.712***
 Treatment (GE)
-0.097
-0.112
---0.055
-0.101
+-0.038
+0.017
+-0.200
+-0.041
 Treatment (CE)
-0.509*
-0.534*
-0.331
-0.730*
+0.333
+0.397
+0.165
+0.502
 Gender (Female)
---0.126
---0.228
+-0.022
+-0.190
+0.105
+-0.377
+Age
+0.227*
+0.196
+0.239*
+0.115
+Education
+-0.043
+-0.079
+-0.090
+0.001
+RecSysUtility
+0.118
+0.051
+0.168
+0.250
+RecSysFrequency
+0.227
+0.230
+0.166
+0.188
+MovieConsumption
+0.099
+0.044
+-0.006
+-0.092
+DisplayOrder
+-0.077
+-0.013
+-0.011
+-0.119
+HistoryNum
+-0.001
+-0.099
+-0.003
+-0.007
+HistoryNum $\times$ MovieConsumption
+0.027
+0.058
+0.007
+0.039
+HistoryNum $\times$ RecSysUtility
+0.009
+-0.040
+-0.021
+-0.006
+Table 6: Mixed-effects linear regression results. Coefficients are reported with standard errors in parentheses. Significance levels: $^{*}p<0.05$, $^{**}p<0.01$, $^{***}p<0.001$.
+C.2 Movie Genre
+Observation of popular movie recommenders like Netflix and Amazon Video showed that most movies recommended tended to fall into widely favored genres such as Comedy and Adventure.
+Based on IMDb genre labels, we annotated each recommended movie in our experimental study as either Comedy or Adventure.
+To examine how genre influences user perceptions and consumption intentions, we performed the following regression analysis:
+$Y_{ij}$ $=\beta_{0}+\beta_{1}\emph{Treatment}_{j}+\beta_{2}\emph{Age}_{j}+\beta_{3}\emph{Gender}_{j}+\beta_{4}\emph{Education}_{j}$
+$\quad+\beta_{5}\emph{MovieConsumption}_{j}+\beta_{6}\emph{RecSysFrequency}_{j}$
+$\quad+\beta_{7}\emph{RecSysUtility}_{j}+\beta_{8}\emph{HistoryNum}_{j}+\beta_{9}\emph{DisplayOrder}_{j}$
+$\quad+\beta_{10}\emph{MvGenre}_{ij}+\beta_{11}\emph{MvGenre}_{j}\times\emph{Treatment}_{j}+u_{0j}+\epsilon_{ij}$
+where ${Y}_{ij}$ denotes the dependent variables (i.e., Cognitive, Affective, Utilitarian, and UserIntent) for the $i$-th observation of the $j$-th subject, with intercepts $\beta_{0}$, subject-level random effects $u_{0j}$, coefficients $\beta_{1}$ through $\beta_{11}$, and an error term $\epsilon_{ij}$.
+Results Regression results (Table 7) show that the genre of the recommended movie significantly influenced user perceptions.
+Specifically, movies categorized as Comedy were associated with lower ratings across affective ($\beta$=-0.203, pp<0.05), utilitarian ($\beta$=-0.149, pp<0.05), and behavioral intention outcomes ($\beta$=-0.415, pp<0.05), compared to Adventure movies.
+The effect of genre was not significantly moderated by explanation type (GE or CE), as the interaction terms were non-significant across all outcomes.
+These results suggest that genre perceptions persist regardless of contextualization strategies and that users may perceive adventure movies as more engaging or informative in recommendation explanations.
+Cognitive Affective Utilitarian UserIntent
+Intercept
+4.946***
+5.045***
+5.189***
+4.879***
+Treatment (GE)
+0.006
+-0.058
+0.034
+-0.268
+Treatment (CE)
+0.438***
+0.135
+0.125
+0.188
+Gender (Female)
+0.116
+0.160
+0.291***
+-0.183
+Age
+-0.075
+-0.089*
+0.013
+-0.148*
+Education
+0.037
+0.028
+0.002
+0.038
+RecSysUtility
+0.306***
+0.298***
+0.273***
+0.297***
+RecSysFrequency
+0.018
+-0.004
+-0.025
+0.130
+MovieConsumption
+0.152**
+0.174***
+0.174***
+0.126*
+DisplayOrder
+0.036
+0.034
+0.016
+0.113**
+HistoryNum
+0.045**
+0.068***
+0.042**
+0.098***
+MvGenre (Com)
+-0.096
+-0.203**
+-0.149*
+-0.415***
+MvGenre (Com) $\times$ Treatment (GE)
+0.014
 0.064
---0.499
-Age
-0.269*
-0.234*
-0.274*
-0.145
-Education
---0.006
---0.040
---0.052
-0.053
-RecSysUtility
-0.217*
-0.155
-0.268
-0.418*
-RecSysFrequency
-0.156
-0.146
-0.076
-0.030
-MovieConsumption
-0.044
---0.018
---0.045
---0.148
-Significance levels: ${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$
-Table 6: Mixed-effects linear regression results for the four dependent variables: Cognitive, Affective, Utilitarian, and UserIntent. Coefficients are shown with standard errors in parentheses.
+0.077
+0.096
+MvGenre (Com) $\times$ Treatment (CE)
+-0.051
+0.068
+0.095
+0.206
+Table 7: Mixed-effects linear regression results. Coefficients are reported with standard errors in parentheses. Significance levels: $^{*}p<0.05$, $^{**}p<0.01$, $^{***}p<0.001$.